# Proxima: a DAG–based cooperative distributed ledger

### v4

@@ 10 Security
where $11/2<\theta<$ 1 is a parameter assumed by Bob.
Note that we can assume that the chain of branches $B_{0},\dots B_{N-1}$ is "longest" in a sense that for $B_{N}=stemPredecesor(B_{N-1})$, $T\notin B_{N}$. In this case:
@@ 10.1 Safety
The central question is under what conditions could an alternative branch $B^{\prime}_{0}$ arrive within the duration of the slot $\tau$, where $\text{slot}(B^{\prime}_{0})=\tau$ and $\text{coverage}(B_{0}^{\prime})>\text{coverage}(B_{0})$, but without containing transaction $T$ (i.e., $T\notin B_{0}^{\prime}$)?
@@
  • In the optimal situation for the attacker, Eve must ensure the amount of "malicious" capital close to the overtaking threshold in every slot. This is not difficult due to fragmentation of the token holdings; See also consideration in case of high concentration of the capital in 10.4
@@ 10.2 Liveness
The confirmation rule introduced in the previous section requires at least $2\cdot\theta\cdot\text{totalSupply}$ of coverage actively contributing to the ledger coverage. If less than $\theta$ of the total on-ledger capital participates, Bob may be waiting indefinitely for the $\beta_{T}(B_{0},\theta)$ to reach the desired value. He will perceive the situation as a liveness problem, meaning his transactions would cease to be confirmed.
@@
One potential approach to alleviating liveness issues is to exclude dormant capital from the calculation of ledger coverage. By introducing additional ledger validity constraints, "sleeping" capital would need to be "awaken", and only after some period of time could it be included in ledger coverage. This could allow for the reduction of the malicious token threshold to a fraction of the active capital.
@@ 10.3 Network partitioning
When participants of the distributed ledger cannot communicate with the rest of the network, we have the situation of dividing (partitioning) the network into several parts. We want to analyze implications of of partitioning to security of the ledger in Proxima.
When a network is partitioned due to communication failures between miners (in PoW), validators (in PoS, BFT), or token holders (in cooperative consensus), each distributed ledger principle behaves differently. We will roughly compare how system behaves in each of the three principles described above. Our intention is to point to the fundamental differences between approaches with respective assumptions and to position Proxima with this regard.
@@ 10.3.1 BFT, PoS
PoS setup (which ultimately always relies on BFT) always assumes a bounded set of validators, a committee. A committee will continue operating as long as the largest partition of the network contains at least $\lfloor 2/3\rfloor+1$ of validators. If this condition is not met, the system stops producing blocks, leading to a liveness issue. However, the system remains secure in the sense that the transactions included in the ledger are considered valid and irreversible343434this is highly simplified model, because many PoS systems employ intricate committee rotation mechanisms which makes liveness analysis not so straightforward. This property is known as deterministic consensus.
@@
In PoW systems such as Bitcoin, the response to network partitions differs significantly from PoS systems. If the miners are isolated into separate partitions, each partition will continue to produce blocks at a reduced rate (due to a lower hashrate in the partition) until the PoW difficulty adjusts accordingly. Consequently, each partition will maintain its own version or fork of the ledger. This situation can persist unnoticed until the network reunites (which will mean that all but one fork will be orphaned and ledger states reverted).
@@
It is important to note that Nakamoto (PoW) systems possess an automatic "self-healing" capability once communication barriers are resolved. Conversely, the process of "healing" in complex BFT/PoS systems may involve more intricate mechanisms.
@@ 10.3.3 Safety and liveness in Proxima
In cooperative consensus, the behavior during network partitioning – specifically among token holders – is akin to that observed in PoW systems.
@@
Upon reconnecting the network partitions and resuming transaction exchange, the nodes will synchronize their UTXO tangles. The biggest ledger coverage rule dictates that sequencers in partitions with a smaller ledger coverage must revert their chains and start anew from the baseline state of the branch with the highest ledger coverage. Forks with lesser coverage will be orphaned, illustrating a form of "self-healing."
@@
In the cooperative consensus, token holders will adopt a similar acceptance rule: waiting until all branches with ledger coverage exceeding $51\%$ of the total token supply include their transaction $T$ (see Section 10.1.Safety).
@@
In summary, Proxima’s cooperative consensus model merges aspects of PoW robustness with novel mechanisms for managing safety and liveness during network partitions, incorporating economic incentives to maintain network integrity.
10.4 Ultimate security. Social consensus
@@
In PoS networks, joining and leaving the committee of validators is not as straightforward, which paradoxically can lead to more decentralized influence compared to completely permissionless PoW networks (although this also depends on other factors, such as how exactly decentralization of the committee is ensured). There are no ultimate decentralization metrics in a BFT/PoS system.
The observation above is also consistent with the demonstration of the social consensus among the dominant Bitcoin miners, who understand that the vast majority of Bitcoin’s real value stems from trust in the network and its expected future. 2-3 largest block proposers (miners) could easily fork the network, for example, with the purpose of changing halving rules. However, they are aware that undermining the trust publicly by compromising decentralization through a protocol fork supported by the majority of hashrate would have unknown consequences, likely resulting in irreversible damage to the Bitcoin narrative and substantial losses for everyone invested in it. Therefore, the fundamental source of Bitcoin’s myth (and, ultimately, its market value) is the strong social consensus among dominant miners. The ledger consensus protocol serves as a tool for coordinating that real-world (social) consensus.
This leads us to another universal observation:
@@
The security of the distributed ledger ultimately depends on the social consensus among entities with the dominant skin-in-the-game.
The cooperative consensus very likely will follow the universal patterns and will converge to a high capital concentration pattern, similar to other fully open-participation networks.
As stated earlier, the Proxima design assumes that capital concentration is inevitable and that the network will be run by a limited number (several dozen, probably up to 100-200) of sequencers, supported by a substantial amount of capital delegated to sequencers by other token holders. In fact, the amount of delegated capital is likely to dominate the capital owned by the sequencers363636we can see strong analogies with how banking sector operates, which normally employs much more leveraged capital than equity in their balance sheets.
@@
11 Spamming prevention
By spamming, we mean all kinds of DDoS-like or storage bloat attack that aims to overwhelm the network and disrupt its operation. We identify several different targets for such attacks and respective countermeasures.
In general, we rely on two main principles for the spamming-prevention: Sybil protection and limit on transaction-per-second rate per user.
The Sybil protection means, that in related attack vectors, the attacker will need to split its holdings among many different fake token holders (addresses) and that will incur cost of storage deposit (see below).
11.1 Spamming the ledger state (state bloat)
Ledger state spamming refers to the act of issuing many small-amount outputs, commonly known as dust. It is also known as the state bloat and is a potential attack vector. The large amount of dust is highly undesirable because it inflates the size of ledger state with a large number of useless UTXOs that never move, because the cost of consumption exceeds the value of locked tokens.
There are well-known strategies to prevent the dusting of the ledger state. The most efficient method is to enforce a minimal number of tokens on each output, based on the size of it in bytes. This is known as a storage deposit. For a token holder, the storage deposit becomes a minimal amount of its holdings. The storage deposit is "refunded" when multiple outputs are compacted into one output.
Proxima adopts the same strategy for dust prevention by enforcing a storage deposit in mandatory output constraints. The token cost per byte is typically a constant in the ledger. However, it may also depend on other deterministic variables, such as the total number of outputs in the ledger state. This allows the cost per byte to increase as the number of outputs approaches certain limits.
The minimal storage deposit may lead to suboptimal consequences in the context of tag-along fees, which may be very small or even zero. Using an ordinary chain-locked output as a tag-along output, a portion of the output is essentially paid as a fee to the sequencer. This can potentially create problems, such as:
@@ 11.1 Spamming the ledger state (state bloat)
  • There is no minimum amount for the output; it can also be a zero-amount output.
@@
  • If it is not consumed within 12 slots, the sender can unlock it for the next 100 slots. This allows the sender to recover any unused fee.
  • If the output is still not consumed, it becomes consumable by any transaction, enabling a "vacuum cleaning" process to remove any remaining dust.
@@ 11.3 Spamming the UTXO tangle
Our approach to prevent spamming the UTXO tangle is to implement rate limits per user (per token holder). This is possible because each transaction (which also is the message between nodes) is signed and identified by the sender’s public key.
@@
Repeated transactions with the same small interval will cause them to be orphaned. If, due to asynchronicity of the network, some transaction will be removed from some nodes and will remain on other nodes, the consistency of the tangle will be restored by the routine solidification mechanism. This is possible because cooperative consensus, being a probabilistic consensus with non-deterministic finality, does not rely on the concepts of transaction being confirmed or rejected.
@@
in general, is a list of accounts or non-fungible assets with their corresponding owners. Ledger state of the UTXO ledger is a set of UTXOs. Past cone of every vertex of the UTXO tangle represents a and a ledger state
@@
is a value of the ledger time stored as a part of the transaction. Timestamp is 00 ticks at genesis. It is said timestamp $\tau$ is on the slot edge when $\tau\mod 128=0$

### v5

@@ 10 Security
where $1/2<\theta<1$ is a parameter assumed by Bob.
Note that we can assume that the chain of branches $B_{0},\dots B_{N-1}$ is "longest" in the sense that for $B_{N}=stemPredecesor(B_{N-1})$, $T\notin B_{N}$. In this case:
@@ 10.1 Safety
The central question of safety is under what conditions could an alternative branch $B^{\prime}_{0}$ arrive within the duration of the slot $\tau$, where $\text{slot}(B^{\prime}_{0})=\tau$ and $\text{coverage}(B_{0}^{\prime})>\text{coverage}(B_{0})$, but without containing transaction $T$ (i.e., $T\notin B_{0}^{\prime}$)?
@@
  • In the optimal situation for the attacker, Eve must ensure the amount of "malicious" capital close to the overtaking threshold in every slot. This is not difficult due to fragmentation of the token holdings; See also consideration in case of high concentration of the capital in 10.6
@@ 10.2 Liveness
The confirmation rule introduced in the previous section requires at least $2\cdot\theta\cdot\text{totalSupply}$ of coverage actively contributing to the ledger coverage. If less than $\theta$ of the total on-ledger capital participates, Bob may be waiting indefinitely for $\beta_{T}(B_{0},\theta)$ to reach the desired value. He will perceive the situation as a liveness problem, meaning his transactions would cease to be confirmed.
@@
10.3 Balancing safety and liveness with $\theta$
The subjective finality rule with parameter $\theta$ (see 10.1) assumed by the user Bob who issues the transaction $T$ means he is waiting for the proof that $T$ was included in the ledgers of other token holders, which together amounts at least $\theta$ part of the totalSupply.
By assuming small values of $\theta$ (close to $1/2$) means Bob faces significant risk, the transaction will be reverted323232that would be equivalent of waiting only for 1 block in the PoW chain.
Taking $\theta$ close to 1, Bob, if lucky, can reach very high safety of his transaction because then $T$ could be reverted only by the consolidated effort of almost all token holders on the network, which is very unlikely.
So, for safety, the larger the parameter $\theta$, the safer the transaction.
On the other hand, large values of $\theta$ means Bob will be waiting until almost all token holders include $T$ in their ledgers. That may never happen and Bob will never convince himself that his transaction $T$ is final.
Therefore, for liveness, the smaller the parameter $\theta$, the easier the network is to perceive as live.
We conclude that the parameter $1/2<\theta<1$ balances the user’s assumptions about safety and liveness. The optimal value of $\theta$ must be somewhere in between the extremes, for example $2/3$ or $3/4$.
It is interesting to compare this finality rule with the longest chain rule in PoW protocols. In the safety part, the rules are somehow similar and both assume an honest majority of 50
However, the liveness aspect is not part of the longest chain rule, because there is no objective proof which part of the hashrate participated in finding the PoW nonce.
@@ 10.4 Network partitioning
When players (peers) in the distributed ledger cannot communicate with the rest of the network, we have the situation of dividing (partitioning) the network into several parts. We want to analyze implications of partitioning for the security of the ledger in Proxima.
When a network is partitioned due to communication failures between miners (in PoW), validators (in PoS, BFT), or token holders (in cooperative consensus), each network behaves differently. We will roughly compare how system behaves in each of the three principles described above. Our intention is to point to the fundamental differences between approaches with respective assumptions and to position Proxima with this regard.
@@ 10.4.1 BFT, PoS
The PoS BFT setup always assumes a bounded set of validators, a committee. A committee will continue operating as long as the largest partition of the network contains at least $\lfloor 2/3\rfloor+1$ of validators. If this condition is not met, the system stops producing blocks, leading to a liveness issue. However, the system remains secure in the sense that the transactions included in the ledger are considered valid and irreversible333333this is highly simplified model, because many PoS systems employ intricate committee rotation mechanisms which makes liveness analysis not so straightforward. This property is known as deterministic consensus.
@@
In PoW systems such as Bitcoin, the response to network partitions differs significantly from BFT systems. If the miners are isolated into separate partitions (for example, for political reasons), each partition will continue to produce blocks at a reduced rate (due to a lower total hashrate in each partition) until the PoW difficulty adjusts accordingly. Consequently, each partition will maintain its own version (fork) of the ledger. This situation can persist unnoticed until the network reunites (which will mean that all but one fork will be orphaned and ledger states reverted).
@@
In the PoW systems, there is no objective way to assess which part of the global hashrate is contributing to the security of the user’s partition. This property is an implication of the anonymity of the players in the consensus.
It is important to note that Nakamoto (PoW) systems possess an automatic "self-healing" capability once communication barriers are resolved. Conversely, the process of "healing" in complex BFT/PoS systems may involve more intricate mechanisms, which require social consensus among the players.
@@ 10.4.3 Safety and liveness in Proxima
In cooperative consensus, the behavior during network partitioning - specifically among token holders – is akin to that observed in PoW systems.
@@
Upon reconnecting the network partitions and resuming the transaction exchange, the nodes will synchronize their UTXO tangles. The biggest ledger coverage rule dictates that sequencers in partitions with a smaller ledger coverage must revert their chains and start anew from the baseline state of the branch with the highest ledger coverage. Forks with lesser coverage will be orphaned, illustrating a form of "self-healing."
@@
In cooperative consensus, token holders will adopt a similar acceptance rule: waiting until all branches with ledger coverage exceeding $51\%$ of the total token supply include their transaction $T$ (see Section 10.1.Safety).
@@
In summary, Proxima’s cooperative consensus model merges aspects of PoW robustness with mechanisms for managing safety and liveness during network partitions, incorporating economic incentives to maintain network integrity.
10.5 Scalability of accounts and liveness
10.5.1 Delegation
The cooperative consensus assumes that a significant majority of tokens must be participating in the consensus all the time. It effectively means that most of tokens must be moved constantly, in each slot.
This fact presents a unique scalability challenge for the protocol, which is absent in other crypto ledgers. We envision at least millions of user accounts that should move their tokens constantly, so how exactly could we do that?
To address this challenge, we introduce the trustless ledger primitives of frozen coverage and delegation. The ledger primitives present special ledger validity constraints (covenants), among many others, and are closely related to the concept of chains.
Here we will describe the delegation covenant of frozen coverage and delegation in general, without going into detail.
The token owner wraps their tokens in the chain output, called the delegation output, and locks it with the delegation lock. The latter allows the specified target sequencer to consume the delegation output and freeze it for a specified but bounded number of slots. For the frozen period of time, only the sequencer can consume the frozen output and only with the purpose of unfreezing it. The frozen funds in the delegation output do not move during the freeze period.
The sequencer is allowed to include the frozen amount in their ledger coverage for the freeze period. In addition, the frozen tokens generate inflation for the sequencer in each slot.
To compensate the delegator for the loan capital, the sequencer must deposit an inflation advance to the delegation output upon freezing it. The protocol of how exactly inflation rewards are shared between the delegator and the sequencer is enforced according to the parameters set by the delegator and the sequencer, as two negotiating sides. In general, the whole process is market driven: for example, if the delegator considers the inflation margin required by the sequencer too high, she moves to the cheaper sequencer (or with better reputation score).
The guarantee that the sequencer cannot steal funds loaned to it is encoded in the transaction validity constraints. After each freeze period, the ledger enforces safe revocation window during which only the delegator can consume the output. So, the delegator is guaranteed full control over her funds after a freeze period.
The maximum freeze period is set by the delegator and can be between 1 and 12 hours. The safe revocation window is around 10 min.
The safe revocation window is just an enforced guarantee for the delegator so that she could always take her funds back. In an overwhelming majority of cases, the sequencer will return funds immediately upon request from the delegator. The request is a safe request on the ledger. Note that in theory the sequencer may ignore the unfreeze request (sequencer is a token holder, a centralized entity). That is expected to be a rare case because by misbehaving, the sequencer will not earn more, yet it will suffer reputation costs.
With the inflation and delegation mechanism, we achieve the following:
  • each token holder is incentivized to delegate their holdings to sequencers of their choice, while enjoying full liquidity of the delegation343434similar to the liquid staking in PoS;
  • sequencer is including coverage of the frozen tokens in each slot without moving the UTXO more often than once per 1 to 12 hours. That results in high liveness and high scalability (with respect to the number of accounts) of the network;
10.5.2 Compulsory freezing
Participation in the consensus is crucial for the safety and liveness of the system. For this reason, all participants in the Proxima ledger are incentivized to participate in the consensus by design.
However, there is no guarantee that all token holders will follow rational behavior for many reasons: neglecting costs, lost private keys, etc.
To address this problem, we introduce special ledger rules that enable and incentivize any sequencer to freeze any UTXO that has not been moved for more than a certain number of slots (say 24 or 48 hours). The compulsory freezing follows the usual delegation rules with the trusted revocation, a maximum of 12 hours of freezing, and the safe revocation window. This guarantees that the owner will never lose her funds, yet the funds will participate in the consensus.
In this way, we ensure high liveness of the system and high participation of the capital in the safety of the ledger without compromising scalability.
10.6 Ultimate security. Social agreement
@@ 10.5.2 Compulsory freezing
In BFT PoS networks, joining and leaving the committee of validators is not as straightforward, which paradoxically can lead to more decentralized influence compared to completely permissionless PoW networks (although this also depends on other factors, such as how exactly decentralization of the committee is ensured). There are no ultimate decentralization metrics in a BFT/PoS system.
The observation above is also consistent with the demonstration of the social agreement among the dominant Bitcoin miners, who understand that the vast majority of Bitcoin’s real value stems from trust in the network and its expected future. 2-3 largest block proposers (miners) could easily fork the network, for example, with the purpose of changing halving rules. However, they are aware that undermining the trust publicly by compromising decentralization through a protocol fork supported by the majority of hashrate would have unknown consequences, likely resulting in irreversible damage to the Bitcoin narrative and substantial losses for everyone invested in it. Therefore, the fundamental source of Bitcoin’s myth (and, ultimately, its market value) is the strong social agreement among dominant miners. The ledger consensus protocol serves as a tool for coordinating that real-world (social) consensus.
This leads us to another universal thesis:
@@
The security of the distributed ledger ultimately depends on the social agreement among entities with the dominant skin in the game.
The cooperative consensus will very likely follow universal patterns and converge to a high capital concentration, similar to other fully open participation networks.
As stated earlier, the Proxima design assumes that capital concentration is inevitable and that the network will ultimately be run by a limited number (probably hundreds, even though the protocol does not set any limits) of sequencers, supported by a substantial amount of capital delegated to sequencers by other token holders. In fact, the amount of delegated capital is likely to dominate the capital owned by the sequencers363636we can see strong analogies with how banking sector operates, which normally employs much more leveraged capital than equity in their balance sheets.
@@
11 Denial-of-service prevention
The denial of service attacks aim to overwhelm the network and disrupt its operation by sending it large amounts of messages. In Proxima those messages are transactions. We identify several different targets for such attacks and respective countermeasures.
In general, we rely on two main principles for the DoS-prevention: Sybil resistance and limit on transaction-per-second rate per user.
The Sybil resistance means, that in related attack vectors, the attacker will need to split its holdings among many different fake token holders (addresses) and that will incur cost of storage deposit (see below).
11.1 State bloat
The ledger state bloat refers to the act of issuing many small-amount outputs, commonly known as dust. The large amount of dust is highly undesirable because it inflates the size of ledger state with a large number of useless UTXOs that never move, because the cost of consumption exceeds the value of locked tokens. It is also a potential attack vector.
Proxima prevents state bloat by enforcing a minimal number of tokens on each UTXO, based on the size of it in bytes. This is known as a storage deposit. For a token holder, the storage deposit becomes a minimal amount of its holdings. The storage deposit is "refunded" when multiple outputs are compacted into one output.
The minimal storage deposit may lead to suboptimal consequences in the context of tag-along fees, which may be very small or even zero. Using an ordinary chain-locked output as a tag-along output, a portion of the output is essentially paid as a fee to the sequencer. This can potentially create problems, such as
@@ 11.1 State bloat
  • There is no requirement for the minimum amount on the tag-along output; it can also be a zero-amount.
@@
  • If it is not consumed within 12 slots, the sender can unlock it during the next 100 slots. This allows the sender to recover any unconsumed tag-along fee.
  • If the output is still not consumed after 100 slots, it becomes consumable by any transaction. This enables a permissionless "vacuum cleaning" process to remove any remaining dust from the ledger.
@@ 11.3 Spamming the UTXO tangle
Our approach to prevent spamming the UTXO tangle is to implement rate limits per user (per token holder). This is possible because each transaction (which is also the message between nodes) is signed and identified by the sender’s public key.
@@
Repeated chain transfers with the small interval will cause them to be orphaned. If, due to asynchronicity of the network, some transaction will be removed from some nodes and will remain on other nodes, the consistency of the tangle will be restored by the routine solidification mechanism. This is possible because cooperative consensus, being a probabilistic consensus with non-deterministic finality, does not rely on the concepts of transaction being confirmed or rejected.
@@
in general, is a list of accounts or non-fungible assets with their corresponding owners. Ledger state of the UTXO ledger is a set of UTXOs (also known as UTXO set). Past cone of every vertex of the UTXO tangle represents a and a ledger state
@@
is a value of the ledger time stored as a part of the transaction. Timestamp is 0 ticks at genesis. It is said timestamp $\tau$ is on the slot edge when $\tau\mod 128=0$